21. A 10 meter tall tank open to the air is filled with a fluid with a density of 1200 kg/m3. It is located with its bottom 10 meters off the ground. When a hole is open at the bottom, fluid sprays out with an initial horizontal velocity. Determine this initial velocity
Answer #1

Using Torricelli's law

v = (2gh)^0.5 = (2*9.8*10)^0.5 = 14 m/s

Which will be required answer.